# Tag-Instruct: Controlled Instruction Complexity Enhancement through Structure-based Augmentation

### v1

@@ Acknowledgements
This project was supported by National Natural Science Foundation of China (No. 62306132) and Guangdong Basic and Applied Basic Research Foundation (No. 2025A1515011564). We thank the anonymous reviewers for their insightful feedbacks on this work.
Appendix A Experimental Setup Details
A.1 Dataset Specifications
@@
A.2 Model Architecture and Training Protocol
@@
For instruction processing, we used Alpaca template for single-turn settings, performed Tag Extraction following methodology from Lu et al. (2023), conducted Tag Reward Estimation using instructions from Xu et al. (2024), and implemented Policy Model Training for tag expansion task.
For additional implementation details, refer to Appendix C.
Appendix B Detailed Description of Baselines
B.1 Methodological Baselines
@@ B.2 Data Baselines
Appendix C Tag Utility Estimation
A. Evaluation Methodology
Our experimental evaluation followed a rigorous iterative process:
Template Selection Process
We selected 1,000 diverse instructions and conducted five iterations of refinement and consolidation to comprehensively evaluate each template variant’s overall utility while maintaining computational efficiency.
Semantic Preservation Measurement
To quantify semantic preservation, we employed two complementary approaches:
Instruction Reconstruction: We utilized specialized encoding and decoding prompts to evaluate the reconstruction quality between original and transformed instructions.
@@ Appendix D Tag Expansion Implementation Details
This section provides comprehensive implementation details for our tag expansion approach. Our method consists of two key components: (1) preference-guided tag sampling and (2) policy optimization.
D.1 Preference-guided Tag Sampling
Given a base tag set $z_{\text{base}}$, we employ a Best-of-N sampling strategy (N=20) to generate candidate tag combinations:
$z_{\text{new}}\sim p(z|z_{\text{base}})$
The sampling process utilizes the tag pool constructed in Section 3.4. For each base tag set, we: 1. Generate N=20 candidate tag combinations 2. Evaluate each candidate using our reward function $R(z_{\text{new}})$ 3. Form preference pairs by selecting: - The highest-scoring combination - One randomly selected combination from the top-k (k=5) candidates
These preference pairs serve as training data for our policy optimization.
D.2 Policy Optimization
We implement two training approaches:
DPO-based Training
The collected preference pairs are used to train a policy through Direct Preference Optimization (DPO). Given a preference pair $(z_{\text{high}},z_{\text{other}})$, we optimize:
@@ D.2 Policy Optimization
$\mathcal{L}_{\text{DPO}}=-\log\sigma(\beta(r_{\theta}(z_{\text{high}})-r_{\theta}(z_{\text{other}})))$
where $r_{\theta}$ is the reward model and $\beta$ is a temperature parameter.
Online Learning
Alternatively, we support online policy learning where the model is updated continuously using the seed instruction data from each iteration. The online approach follows: 1. Initialize policy $\pi_{\theta}$ with pretrained weights 2. For each iteration $t$: - Sample tag combinations using current policy - Collect rewards and update policy using PPO - Update seed instruction pool with generated examples
D.3 Implementation Details
Key hyperparameters and architectural choices: - Sampling temperature: 0.7 - PPO clip ratio: 0.2 - KL penalty coefficient ($\beta$): 0.1 - Learning rate: 5e-5 - Batch size: 64 preference pairs - Training iterations: 1000 per phase
The reward components are weighted as follows: - $\alpha_{1}$ (utility): 0.4 - $\alpha_{2}$ (alignment): 0.3 - $\alpha_{3}$ (coherence): 0.3
Appendix E Encode/Decode Prompts for I2T and T2I
Appendix F Example Appendix
Tag Combination Prompt Template Below is an instruction that describes a task, write a response that appropriately completes the request. ### Instruction: Create a comprehensive and challenging task that incorporates these concepts: {tags}. Do not explain or provide any additional information. Only return the task/instruction. ### Response:
Figure 5: Tag Expansion and Complexity Enhancement Template
Appendix G Dataset Attributes
Appendix H Tag Extraction Experiment
@@
We selected 1,000 instructions from alpaca-clean  Taori et al. (2023) and conducted five iterations of tag expansion to obtain a total of 5k instructions. We then followed the same supervised fine-tuning process as in the previous section. Our aim was to evaluate the performance of different encoding strategies in generating tags from instructions.
To systematically evaluate different encoding approaches, we designed four template variants that explore different aspects of instruction encoding:
@@
Each template was carefully designed to test specific hypotheses about effective instruction encoding, with variations in guidance specificity, semantic preservation requirements, and intent extraction mechanisms (detailed templates in Appendix H).
@@
We present four different template designs for tag generation and analyze their performance:
@@
Comprehensive Tag Extraction Prompt Template (Reconstructed Rate: 0.5373  Arena Hard: 0.1992) You are a semantic analysis expert. Your task is to extract exactly three essential tags that capture the core semantic concepts of the given instruction or question. The tags should be: - Concise (1–2 words each) - Hierarchically ordered by importance - Generalizable across similar tasks ### Guidelines: 1. Focus on action-oriented concepts 2. Avoid redundant or overlapping tags 3. Use standard terminology when possible ### Examples: [Input] In the context of climate change adaptation, analyze how urban planning strategies can be modified to create resilient cities. [Tags] urban resilience, climate adaptation, sustainable development [Input] Design a comprehensive employee training program for a multinational corporation addressing cultural sensitivity. [Tags] corporate training, global workforce, skill development ### Task: Given the following instruction, provide exactly three semantic tags following the above format and guidelines: [Input] {instruction} [Output Format] tag1, tag2, tag3 [Tags]
Explanatory Tag Annotation Prompt Template (Reconstructed Rate: 0.36  Arena Hard: 0.1667) You are a helpful assistant. Please identify tags of user intentions in the following user query and provide an explanation for each tag. Please respond in the JSON format: { "tag": "str", "explanation": "str" } Query: {instruction} Assistant:
@@
Appendix I Experimental Setup
I.1 Tag Analysis Prompt
@@ 1. A method of delivering computing services and storing data remotely over the internet
I.2 Tag-based Control Analysis.
Prompt Template for Self-Instruct Task Generation Generate a query based on the following query. ### Important: 1. The query must differ from the examples provided. 2. Don’t provide a solution or answer to the query. 3. Output Format is as follows: Output: [query]: xxx ### Example: [query]: {query1} [query]: {query2} [query]: {query3} [query]: {query4} Output:
Prompt Template for Tag-Based Instruction Generation Generate new tags and a corresponding query based on the following tags and query. ### Important: 1. The new tags and query must differ from the examples provided. 2. Don’t provide a solution or answer to the query. 3. Output Format is as follows: Output: [tags]: xxxx, xxx, xx [question]: xxx ### Example: [tags]: {tag1} [question]: {question1} [tags]: {tag2} [question]: {question2} [tags]: {tag3} [question]: {question3} [tags]: {tag4} [question]: {question4} Output:
Appendix J Tag Analysis Results

### v2

@@ Acknowledgements
This project was supported by National Natural Science Foundation of China (No. 62306132) and Guangdong Basic and Applied Basic Research Foundation (No. 2025A1515011564). The paper was also supported by the Key Project of the Shanghai Municipal Education Commission’s AI-Enabled Research Paradigm Reform and Discipline Leap Program (Development of a Domain-Specific Large Language Model in the Field of Urban and Rural Planning for Enhancing Spatial Cognition and Decision-Making Capabilities) and by the Fundamental Research Funds for the Central Universities (22120250239). This work was done by He Zhu during his internship at SUSTech. We thank the anonymous reviewers for their insightful feedbacks on this work.
Appendix A Tag-Instruct Prompt Template Demonstration
A.1 Encode/Decode Prompts for I2T and T2I
Figure 5: Prompt template for extracting semantic tags from instructions, with examples demonstrating tag extraction across diverse domains
Figure 6: Template for generating enhanced instructions from semantic tags while maintaining clarity and solvability
Figure 7: Template for expanding tag set and enhancing question complexity through contextual tag addition
Figure 8: Simple template for generating responses to task instructions
Appendix B Experimental Setup Details
B.1 Dataset Specifications
@@
B.2 Model Architecture and Training Protocol
@@
For instruction processing, we used the Alpaca template for single-turn settings, performed tag extraction following methodology from Lu et al. (2023), conducted tag reward estimation using instructions from Xu et al. (2024), and implemented policy model training for the tag expansion task. For additional implementation details, refer to Appendix D.
Appendix C Detailed Description of Baselines
C.1 Methodological Baselines
@@ Appendix D Tag Expansion Implementation Details
This section provides implementation details for our tag expansion approach using reinforcement learning to optimize the policy model for high-utility tag generation.
D.1 Tag Pool Construction
We construct a comprehensive tag pool using Magpie-500k Xu et al. (2024) as seed data:
Tag Extraction and Utility Estimation
For each instruction in Magpie-500k, we extract semantic tags using the teacher LLM following Section 3.2. We then compute the utility score for each unique tag using Equation 1:
$\phi_{i}=\frac{\sum_{j\in\mathcal{D}_{i}}|y_{j}|}{|\mathcal{D}_{i}|}$
where $\mathcal{D}_{i}$ denotes instructions containing tag $i$.
Tag Pool Formation
Based on utility scores, we rank all extracted tags and form the tag pool $\mathcal{T}=\{t_{1},t_{2},\ldots,t_{N}\}$ ordered by decreasing utility. We define $\mathcal{T}_{\text{good}}$ as the top 10,000 highest-utility tags and $\mathcal{T}_{\text{bad}}$ as the bottom 10,000 lowest-utility tags.
D.2 Best-of-N Tag Expansion and Scoring
Candidate Generation
For each base tag set $z_{\text{base}}$, we follow the prompt template in Figure 7 to generate $N=20$ candidate expansion tags using the teacher LLM:
$z_{\text{new}}^{(i)}\sim p(z|z_{\text{base}},x),\quad i=1,2,\ldots,20$
Embedding-based Scoring
We score each candidate tag $z_{\text{new}}^{(i)}$ using embedding similarity with the pre-constructed tag pools. Let $\text{emb}(\cdot)$ denote the embedding function, we compute:
$\text{score}_{\text{good}}^{(i)}$ $=\frac{1}{|\mathcal{T}_{\text{good}}|}\sum_{t\in\mathcal{T}_{\text{good}}}\cos(\text{emb}(z_{\text{new}}^{(i)}),\text{emb}(t))$ (2)
$\text{score}_{\text{bad}}^{(i)}$ $=\frac{1}{|\mathcal{T}_{\text{bad}}|}\sum_{t\in\mathcal{T}_{\text{bad}}}\cos(\text{emb}(z_{\text{new}}^{(i)}),\text{emb}(t))$ (3)
The final score for candidate $i$ is:
$\text{score}^{(i)}=\text{score}_{\text{good}}^{(i)}-\text{score}_{\text{bad}}^{(i)}$
The insight is that high-quality tags should be semantically closer to high-utility tags and farther from low-utility tags.
D.3 Preference Pair Construction
Based on the computed scores, we construct preference pairs for DPO training. For the chosen tag, we select $z_{\text{chosen}}=\arg\max_{i}\text{score}^{(i)}$ which represents the highest-scoring candidate. For the rejected tag, we randomly sample $z_{\text{rejected}}$ from the remaining candidates $\{z_{\text{new}}^{(j)}:j\neq\arg\max_{i}\text{score}^{(i)}\}$.
D.4 Policy Model Training
DPO Training
The policy model is optimized using Direct Preference Optimization on the collected preference pairs:
@@ D.4 Policy Model Training
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(z_{w},z_{l})}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(z_{w}|z_{\text{base}},x)}{\pi_{\text{ref}}(z_{w}|z_{\text{base}},x)}-\beta\log\frac{\pi_{\theta}(z_{l}|z_{\text{base}},x)}{\pi_{\text{ref}}(z_{l}|z_{\text{base}},x)}\right)\right]$
where $z_{w}$ and $z_{l}$ denote chosen and rejected tags respectively, $\pi_{\text{ref}}$ is the reference model (teacher LLM), and $\beta$ controls the KL penalty strength. We trained the model using a learning rate of $5\times 10^{-5}$ with batch size of 128 preference pairs over 1 epochs. The KL penalty coefficient ($\beta$) was set to 0.1, and we used the sentence-transformers/all-MiniLM-L6-v2 model for generating embeddings.
While the above process can also be implemented using Proximal Policy Optimization (PPO) with the reward function:
$R(z_{\text{new}})=\text{score}_{\text{good}}-\text{score}_{\text{bad}}$
and the PPO objective:
$\mathcal{L}_{\text{PPO}}=\mathbb{E}_{t}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right]$
where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\text{old}}(a_{t}|s_{t})}$ and $\hat{A}_{t}$ is the advantage estimate, we choose DPO for its superior efficiency and stability in practice.
Appendix E Analysis Experiment Setup
E.1 Tag Extraction Experiment
@@
We selected 1,000 instructions from alpaca-clean  Taori et al. (2023) and conducted five iterations of tag expansion to obtain a total of 5k instructions. We then followed the same supervised fine-tuning process as in the previous section. Our aim was to evaluate the performance of different encoding strategies in generating tags from instructions. To systematically evaluate different encoding approaches, we designed four template variants that explore different aspects of instruction encoding:
@@
Each template was carefully designed to test specific hypotheses about effective instruction encoding, with variations in guidance specificity, semantic preservation requirements, and intent extraction mechanisms (detailed templates in Appendix E.1).
@@
We present three different template designs for tag generation and analyze their performance, with the Evolved template being our I2T prompt shown in Figure 5:
@@
Enhanced Tag Generation Prompt Template (Reconstructed Rate: 0.5373  Arena Hard: 0.1992) You are a semantic analysis expert. Your task is to extract exactly three essential tags that capture the core semantic concepts of the given instruction or question. The tags should be: - Concise (1–2 words each) - Hierarchically ordered by importance - Generalizable across similar tasks ### Guidelines: 1. Focus on action-oriented concepts 2. Avoid redundant or overlapping tags 3. Use standard terminology when possible ### Examples: [Input] In the context of climate change adaptation, analyze how urban planning strategies can be modified to create resilient cities. [Tags] urban resilience, climate adaptation, sustainable development [Input] Design a comprehensive employee training program for a multinational corporation addressing cultural sensitivity. [Tags] corporate training, global workforce, skill development ### Task: Given the following instruction, provide exactly three semantic tags following the above format and guidelines: [Input] {instruction} [Output Format] tag1, tag2, tag3 [Tags]
Model-based Tag Annotation Prompt Template (Reconstructed Rate: 0.36  Arena Hard: 0.1667) You are a helpful assistant. Please identify tags of user intentions in the following user query and provide an explanation for each tag. Please respond in the JSON format: { "tag": "str", "explanation": "str" } Query: {instruction} Assistant:
@@
E.2 Derived Meaning Prompt
@@ 1. A method of delivering computing services and storing data remotely over the internet
Figure 10: Template for analyzing tag semantics through divergent thinking and convergent analysis.
E.3 Tag Combination
Building on our quantitative control framework through high-utility tags, we introduce tag combination as a principled approach to instruction synthesis. Tag combination merges multiple semantic concepts to create richer, more complex instructions. While prior work has explored combining concepts heuristically, our utility-based selection provides a systematic way to identify and combine high-value tags for optimal instruction generation. The following template (see Figure 11) demonstrates our approach to combining tags into coherent instructions.
Tag Combination Prompt Template Create a comprehensive and challenging task that incorporates these concepts: {tags}. Do not explain or provide any additional information. Only return the task/instruction. ### Response:
Figure 11: Template for combining tags into a task that incorporates their semantic concepts.
Appendix F Case Study: Tag Examples
@@
Appendix G Case Study: Progressive Enhancement of Web Development Requirements
This case study demonstrates the progressive enhancement of web development requirements, showing how a simple book listing page evolves into a comprehensive, accessible, and feature-rich web application. The following table presents five iterations of requirements, each building upon the previous version.
Version Requirements Tags
Iter1 Develop an HTML page that dynamically displays a list of books and their authors using CSS for styling. The page should include a search functionality that filters the book list based on the author’s name. Additionally, implement a hover effect that reveals a brief description of each book when the user hovers over a book title. Ensure the page is responsive and visually appealing. web_development, book_listing, author_info, css_styling
Iter2 Develop an HTML page that dynamically displays a list of books and their authors using CSS for styling. The page should include a search functionality that filters the book list based on the author’s name, as well as a pagination system to handle a large number of books. Additionally, implement a hover effect that reveals a brief description of each book when the user hovers over a book title. Ensure the page is responsive and visually appealing, and optimize the user experience by incorporating smooth transitions and animations for the hover effect and pagination. web_development, user_interface, responsive_design, user_experience
Iter3 Develop an HTML page that dynamically displays a list of books and their authors using CSS for styling. The page should include a search functionality that filters the book list based on the author’s name, as well as a pagination system to handle a large number of books. Additionally, implement a hover effect that reveals a brief description of each book when the user hovers over a book title. Ensure the page is responsive and visually appealing, and optimize the user experience by incorporating smooth transitions and animations for the hover effect and pagination. Furthermore, enhance the accessibility of the page by ensuring that all interactive elements are keyboard navigable and that the page is compatible with screen readers. Consider the use of ARIA roles and properties to improve the accessibility of the search functionality and pagination system. web_development, user_interface, responsive_design, accessibility
Iter4 Develop an HTML page that dynamically displays a list of books and their authors using CSS for styling. The page should include a search functionality that filters the book list based on the author’s name, as well as a pagination system to handle a large number of books. Additionally, implement a hover effect that reveals a brief description of each book when the user hovers over a book title. Ensure the page is responsive and visually appealing, and optimize the user experience by incorporating smooth transitions and animations for the hover effect and pagination. Furthermore, enhance the accessibility of the page by ensuring that all interactive elements are keyboard navigable and that the page is compatible with screen readers. Consider the use of ARIA roles and properties to improve the accessibility of the search functionality and pagination system. Additionally, implement a feature that allows users to sort the book list by title, author, or publication year. Ensure that the sorting functionality is accessible and provides clear visual feedback to users. Finally, optimize the page’s performance by implementing lazy loading for images and minimizing the use of heavy CSS animations to ensure a smooth user experience on devices with limited processing power. web_development, user_experience, accessibility, performance_optimization
Table 7: Case Study of Progressive Enhancement of Web Development Requirements